OpenStudy (anonymous):

If the simple interest on $5000 for 9 years is $3600, then what is the rate?

OpenStudy (anonymous):

A= P(1+ rt)

OpenStudy (anonymous):

3600= 5000(1+(r) (9))

OpenStudy (anonymous):

3600= 5000 + 9r 3600= 5000 + 45000r -5000 -5000 -1400 = 45000r